A hockey puck with an initial speed of 5.0 m/s coasts 50 m to a stop across the ice. If the force of friction on the puck has a magnitude 0.12 N, what is the puck’s mass?
1
Expert's answer
2020-11-30T15:40:20-0500

Answer

Using energy conservation

work done = kinetic energy

"Fd=\frac{mv^2}{2}"

Mass is given by

"m=\frac{2Fd}{v^2}=\frac{2\times0.12\times50}{25}"


m=0.48Kg